OpenStudy (anonymous):

Medal Solve by the substitution method 5x + 3y=8 3x +y = 8

pooja195 (pooja195):

3x +y = 8

pooja195 (pooja195):

What can we do so that the Y will be alone?

OpenStudy (anonymous):

move it to the right side

OpenStudy (anonymous):

or move the 3x to the other side

pooja195 (pooja195):

good y=-3x+8 whats next?

OpenStudy (anonymous):

sub in the other equation

pooja195 (pooja195):

good :) \[\huge~5x + 3(-3x+8)=8 \]

pooja195 (pooja195):

Solve that

OpenStudy (anonymous):

5x -9x + 24 = 8

pooja195 (pooja195):

now combine the like terms :)

OpenStudy (anonymous):

5x - 9x = 8 -24

pooja195 (pooja195):

ok 8-24=?

OpenStudy (anonymous):

-4x / -4x = -16 / -4

OpenStudy (anonymous):

answer is 4

OpenStudy (anonymous):

x = 4

pooja195 (pooja195):

Good now what?

OpenStudy (anonymous):

we sub in 4 to the other equation

pooja195 (pooja195):

good \[\huge~y=-3(4)+8 \]

pooja195 (pooja195):

Solve

OpenStudy (anonymous):

y = -12x + 8

pooja195 (pooja195):

Where did u get x from?

pooja195 (pooja195):

Remeber we substituted 4 in place of x

OpenStudy (anonymous):

ohh yeah sorry

OpenStudy (anonymous):

so what is the next step

pooja195 (pooja195):

Solve :) y=-3(4)+8 -3 x 4 = ?

OpenStudy (anonymous):

-12

pooja195 (pooja195):

-12 + 8 = ?

OpenStudy (anonymous):

y = -12 + 8

pooja195 (pooja195):

-12+8=?

OpenStudy (anonymous):

-4

pooja195 (pooja195):

good :) so x= ? y=?

OpenStudy (anonymous):

x= 4 and y= -4

pooja195 (pooja195):

yep :)

pooja195 (pooja195):

(4,-4)

OpenStudy (anonymous):

so is this the answer?

pooja195 (pooja195):

yes

OpenStudy (anonymous):

thank you soo much :)

pooja195 (pooja195):

:)